OpenStudy (anonymous):

please help Determine the solution set of the system. x^2 + y^2 = 49 x + y = 7

OpenStudy (hoblos):

solve the second equation for y replace the result in the first equation you will have a quadratic equation do you know how to solve quadratic equations ?

OpenStudy (anonymous):

nevermind. can u just tell me how did they get ''-14'' from this? x^2+(7-x)^2=49 x^2+49-14x+x^2=49

OpenStudy (phi):

first add -49 to both sides. what do you get ?

OpenStudy (solomonzelman):

circle and a line ;)

OpenStudy (hoblos):

(a-b)^2 = a^2 -2ab + b^2 so (7-x)^2 = 49-14x+x^2

OpenStudy (anonymous):

yeah...but how did u get -14x

OpenStudy (hoblos):

follow the formula (a-b)^2 = a^2 -2ab + b^2 (7-x)^2 = 7^2 -2(7)(x) + x^2 = 49-14x+x^2

OpenStudy (anonymous):

ohh okay thanks guys:)

OpenStudy (hoblos):

any time :)